# Statistical features of quantum chaos using the Krylov operator complexity

### v2

@@ 4 Model and Results
In Figure 5, the average correlation matrix $\left\langle x_{i}x_{j}\right\rangle$ is shown for the case of GUE. With $\mathcal{N}_{max}=50$, its histogram approaches the normal distribution as shown in Figure 1(b) so that there is enough randomness in the statistics. There is a clear distinction in the pattern between the chaotic and the nonchaotic case. In the nonchaotic case, there are heavy cross-like structures with large matrix values. But in the chaotic case, these heavy cross-like structure disappears (or fades away) and matrix values are more uniformly or smoothly distributed.
In Figure 6, the variances $\sigma^{2}$ are shown for the samples of different distributions (GOE, GUE, URE, UIM and UCP) of initial operators, where $m$ labels the sample. Note that only the first 1000 out of the total 5000 samples are shown for the purpose of visual clearness, and the remaining samples have similar behaviors. For GOE and GUE, they overlap completely in the nonchaotic case in Figure 6(b), and separate from each other in the chaotic case in Figure 6(a). Similary for the uniform distributions, they mix together in the nonchaotic case in Figure 6(d), and split into two separated groups in the chaotic case in Figure 6(c), with URE and UIM being one group and UCP being the other.
In Figure 7, the histogram of all the sampled $\sigma^{2}$ is shown and the resulting probability distribution resembles the normal distribution. For the Gaussian ensembles, we can see clearly that they are overlapped in the nonchaotic case in Figure 7(b) and separated in the chaotic case in Figure 7(a). For the uniform distributions, they are overlapped in the nonchaotic case in Figure 7(d). In the chaotic case in Figure 7(c), URE and UIM are still together, but they are separated from UCP.
Figures 6 and  7 straightforwardly show that the resulting distributions $f_{\sigma^{2}}$ can distinguish the chaotic and the nonchaotic behaviors. Their axis values indicate that they can be compared together in one plot. So we fit 333 The details of this data fitting is shown in the Appendix A. their histogram to the normal distribution and find the average value $\langle\sigma^{2}\rangle$ and the standard deviation, which can characterize the resulting distribution $f_{\sigma^{2}}$. Then the resulting distribution $f_{\sigma^{2}}$ can be shown in one plot where the point is this average value and the error bar is the standard deviation. In Figure 8, the resulting distributions $f_{\sigma^{2}}$ are shown for all different distributions (GOE, GUE, URE, UIM and UCP) of initial operators, as the system varies from nonchaotic $a=0$ to chaotic $a=1$. When the system is nonchaotic, the data points of all distributions overlap together within the range of error bars. As the system becomes chaotic, different distributions start to separate approximately at $a=0.1$ and the distance of separation is roughly fixed after $a=0.15$. Note that the axis is logarithmic, so the separation is very large. We see that they split into two groups well-separated: one group is of GOE, URE and UIM, the other group is GUE and UCP.
5 Discussion
@@
To begin, the average correlation matrix $\left\langle x_{i}x_{j}\right\rangle$ has different patterns between the nonchaotic and the chaotic case, as shown in Figure 5. From the aspect of RMT, the average correlation matrix $\left\langle x_{i}x_{j}\right\rangle$ is the common estimator of the statistics. So different patterns qualitatively tell us that the statistics of $\left\langle x_{i}x_{j}\right\rangle$ can distinguish between the nonchaotic and the chaotic case. On the other hand, $x_{i}$ characterizes the disorderedness of Lanczos coefficients from the aspect of the Krylov chain. So $\left\langle x_{i}x_{j}\right\rangle$ quantitatively captures the erratic behavior and the correlation of Lanczos coefficients. In the nonchaotic case of Figure 5(b), the large matrix values indicate more erratic behavior of Lanczos coefficients and more localization on the Krylov chain. So it provides another view for the connection between chaos and Anderson localization and is consistent with results of Dymarsky:2019elm ; Rabinovici:2021qqt .
@@
The repetition of the heavy cross-like structure in Figure 5(b) is also a very interesting phenomenon. It might come from some abstract periodic boundary conditions of the Krylov chain. Or it might be connected to the phenomenon of two-electron correlations of Anderson localization on the Krylov chain Weinmann1995h2eOF . In Figure 9, we plot the average correlation matrix of the wave function $\langle\ln{|\psi_{2m}\psi_{2n}|}\rangle$ for the zero frequency state of the Krylov chain. There are also heavy cross-like structures for the nonchaotic case in Figure 9(b). These cross-like structures locally look like the cross-like structure of the two-electron correlations of Anderson localization in disordered mesoscopic ring Weinmann1995h2eOF . And the cross-like structure of $\ln{|\psi_{2m}\psi_{2n}|}$ disappears in the chaotic case in Figure 9(a). The physics of this structure remains an open problem at this stage.
Then, we emphasize that these overlapping-separating behaviors do not depend on the choice of algorithm step-sizes of coefficients $\left\{b_{n}\right\}$. In appendix B, we choose the step-sizes $10\mathcal{N}_{max}\leq n\leq 15\mathcal{N}_{max}$ from the Krylov space in Figure 4. We still get the resulting normal distributions and the overlapping-splitting behaviors. Note that the detailed values of the average and the standard deviation will change, it is the overlapping-separating behavior that does not change. It remains an open problem to quantitatively prove this independence on different choices of algorithm step-sizes.
In addition, we emphasize the importance of the resulting distribution being the normal distribution. We repeat the computation for $\mathcal{N}_{max}=5$ in Appendix C, where the resulting distribution $f_{\sigma^{2}}$ is not normal and there is no separation of $f_{\sigma^{2}}$ in the chaotic case. This suggests that the resulting distributions being normal is essential for their separation behavior in the chaotic case.
Qualitatively we can understand this by the central limit theorem as discussed in Section 3. The resulting distribution being normal means that we have enough randomness in the sample, so that the statistics is reliable and robust. Empirically, a choice of $\mathcal{N}_{max}=50$ is already adequate to get the normal distribution. It remains an open problem to quantitatively understand the connection between being normal distributions and the separation behavior of $f_{\sigma^{2}}$.
@@
After that, we discuss the resulting two groups of separation. The GOE, URE and UIM are in one group. We can view GOE, URE and UIM as essentially composed of only real numbers, because the UIM is just real numbers multiplied by the overall factor-the imaginary $i$. The other group is GUE and UCP, which are essentially composed of complex numbers with both real and imaginary parts. So one group is essentially related with one-dimensional random numbers and the other group is related with two-dimensional random numbers. This distinction between real and complex numbers may be related to the nature of the integrability-breaking-term, or to the symmetry of the model haake1991quantum . This remains to be one of the open problems.
@@
In conclusion, we find two statistical quantities that have distinct behaviors as the Sinai billiard system changes from nonchaotic to chaotic, i.e., the average correlation matrix $\langle x_{i}x_{j}\rangle$ and the resulting distribution $f_{\sigma^{2}}$ of the variance of Lanczos coefficients. The statistics of these two quantities are the Wishart distribution and the rescaled chi-square distribution respectively, independent of distributions of initial operators. When the chi-square distribution approaches the normal distribution, the resulting distributions $f_{\sigma^{2}}$ can distinguish the nonchaotic and the chaotic regime: they are overlapped in the nonchaotic case and they split into two groups in the chaotic case. The average correlation matrix $\langle x_{i}x_{j}\rangle$ also exhibits different patterns in the nonchaotic and the chaotic regime: there are heavy cross-like structures in the nonchaotic case. These results provide a dual view on results in the literature and reveal a deeper connection among Lanczos coefficients, RMT and the Anderson localization.
@@
(a) $\left|\langle x_{i}x_{j}\rangle\right|$ in chaotic case $a=1$.
(b) $\left|\langle x_{i}x_{j}\rangle\right|$ in nonchaotic case $a=0$.
Figure 5: Patterns of $\left|\langle x_{i}x_{j}\rangle\right|$ with $\mathcal{N}_{max}=50$
@@
[t] Refer to caption
Figure 8: The logarithm plot for the average $\langle\sigma^{2}\rangle$ of the resulting normal distribution $f_{\sigma^{2}}$, for all different distributions of initial operators. The error bar is the standard deviation. The system dynamics changes from nonchaotic to chaotic, as $a$ varies from 00 to $1$.
(a) $\left\langle\ln{|\psi_{2m}\psi_{2n}|}\right\rangle$ in chaotic case $a=1$.
(b) $\left\langle\ln{|\psi_{2m}\psi_{2n}|}\right\rangle$ in nonchaotic case $a=0$.
Figure 9: Patterns of $\left\langle\ln{|\psi_{2m}\psi_{2n}|}\right\rangle$ with $\mathcal{N}_{max}=50$
Acknowledgements.
Appendix A Check the normal distribution of $\sigma^{2}$ for $\mathcal{N}_{max}=100$
Here we show the data fitting of the histogram in Figure 7 to the normal distribution. For GOE and GUE, the fitted distribution is shown in Figure 10. For uniform distributions, the fitted distribution is shown in Figure 11. The average value and the standard deviation are listed in Table 1, where we compare the values obtained from the sampled data and from the fitted normal distribution.
Appendix B Different choice of algorithm step sizes for $b_{n}$
For the overlapping-separating behavior, here we show its independence on the sets of coefficients used in computing $\sigma^{2}$. Now we choose the Lanczos coefficients $\left\{b_{n}\right\}$ from step-sizes $10\mathcal{N}_{max}\leq n\leq 15\mathcal{N}_{max}$ of Figure 4.
Firstly, we compute the variance $\sigma^{2}$ (3.3), where the first $1000$ samples are shown in Figure 12. We see that they mix together in the nonchaotic case and separate from each other in the chaotic case.
Then, the histogram of all $5000$ samples are shown in Figure 13. Obviously they still resemble normal distributions and are overlapped in the nonchaotic case and well-separated in the chaotic case.
Finally, the fitted normal distributions are shown in Figure 14 and 15. So we still get the resulting normal distributions and the overlapping-splitting behaviors, although the detailed values of the average and the standard deviation change.
Appendix C Check the chi-square distribution of $\sigma^{2}$ for $\mathcal{N}_{max}=5$
Here we study the case of $\mathcal{N}_{max}=5$. Then the sets of coefficients $b_{n}$ are chosen from step-sizes $151<n<$ 15. We repeat the computation and show that there is neither normal distributions nor separating behaviors.
The histogram of all $5000$ samples are shown in Figure 16. Obviously they are not normal distributions and are not separated in the chaotic case. So the normal distribution is necessary for the overlapping-separating behavior.
We fit the histogram to the rescaled chi-square distribution, the results are shown in Figure 17 for the Gaussian ensembles and in Figure 18 for the uniform distributions,. We see that the rescaled chi-square distribution indeed captures the resulting distribution $f_{\sigma^{2}}$, so it verifies our analysis in Section 3.3.
@@ Appendix C Check the chi-square distribution of $\sigma^{2}$ for $\mathcal{N}_{max}=5$
(a) Histogram of $\sigma^{2}$ of GOE and GUE in the chaotic case $a=1$.
(b) Histogram of $\sigma^{2}$ of GOE and GUE in the nonchaotic case $a=0$.
(a) $N=5$ chaos, GUE. Skewness=1.19843, Kurtosis=5.18562.
(b) $N=5$ nonchaos, GUE. Skewness=4.30801, Kurtosis=50.7359.
(c) $N=5$ chaos, GOE. Skewness=1.6231, Kurtosis=7.05429.
(d) $N=5$ nonchaos, GOE. Skewness=5.83862, Kurtosis=73.6201.
(a) Chaotic case $a=1$ of UIM. Skewness=0.306957, Kurtosis=3.26486.
(b) Chaotic case $a=1$ of URE. Skewness=0.234259, Kurtosis=2.88038.
(c) Chaotic case $a=1$ of UCP. Skewness=0.207669, Kurtosis=2.97609.
x

### v3

@@ 4.2 Behaviors between the nonchaotic and the chaotic case
[I] Correlation matrix: For the Sinai billiard, the average correlation matrix $\left\langle x_{i}x_{j}\right\rangle$ has different patterns between the nonchaotic and the chaotic case, shown in Figure 4(a) and  4(b) for the case of GUE. With $\mathcal{N}_{max}=50$, its histogram approaches the normal distribution as shown in Figure 2(b) so that there is enough randomness in the statistics. There is a clear distinction in the pattern between the chaotic and the nonchaotic case. In the nonchaotic case, there are heavy cross-like structures with large matrix values. But in the chaotic case, these heavy cross-like structure disappears (or fades away) and matrix values are more uniformly or smoothly distributed.
For the Stadium billiard, the average correlation matrix $\left\langle x_{i}x_{j}\right\rangle$ has no clear distinction between the chaotic and the nonchaotic case, as shown in Figure 4(c) and  4(d) for the case of GUE. With $\mathcal{N}_{max}=50$, its histogram approaches the normal distribution as shown in Figure 2(d) so that there is enough randomness in the statistics. However, matrix values are smoothly distributed without cross-like structures in both the chaotic and the nonchaotic case.
From the aspect of RMT, the average correlation matrix $\left\langle x_{i}x_{j}\right\rangle$ is the common estimator of the statistics. The similarity of Figure 4(a),  4(c) and  4(d) tells us that they have similar statistics. This is interesting because the Stadium billiard in all phases, whether chaotic or not, are close to the nonchaotic Sinai billiard.
@@
On the other hand, $x_{i}$ characterizes the disorderedness of Lanczos coefficients from the aspect of the Krylov chain. So $\left\langle x_{i}x_{j}\right\rangle$ quantitatively captures the erratic behavior and the correlation of Lanczos coefficients. The large matrix values for the nonchaotic Sinai billiard (Figure 4(b)) indicate more erratic behavior of Lanczos coefficients and more localization on the Krylov chain. The Stadium billiard (Figure 4(c),  4(d)) and the chaotic Sinai billiard (Figure 4(a)) mean that less erratic behavior there.
@@
Furthermore, the repetition of the heavy cross-like structure in the nonchaotic Sinai billiard (Figure 4(b)) is a very interesting phenomenon. It might come from some abstract periodic boundary conditions of the Krylov chain. Or it might be connected to the phenomenon of two-electron correlations of Anderson localization on the Krylov chain Weinmann1995h2eOF . To make this connection with Anderson localization, we plot the average correlation matrix of the wave function $\langle\ln{|\psi_{2m}\psi_{2n}|}\rangle$ for the zero frequency state of the Krylov chain. For the nonchaotic Sinai billiard, there are also heavy cross-like structures for the correlation matrix of the wave function as shown in Figure 5(c). These cross-like structures locally look like the cross-like structure of the two-electron correlations of Anderson localization in disordered mesoscopic ring Weinmann1995h2eOF . And the cross-like structure of $\ln{|\psi_{2m}\psi_{2n}|}$ disappears in the chaotic Sinai billiard in Figure 5(b). For the Stadium billiard, the correlation matrix of the wave function is shown in Figure 6(b) for the chaotic case and in Figure 6(c) for the nonchaotic case. They are nearly the same and is also similar to the chaotic Sinai billiard (Figure 5(b)), without cross-like structure. This behavior is consistent with the statistics $\left\langle x_{i}x_{j}\right\rangle$.
[I] distribution of the variance $\sigma^{2}$: For the Sinai billiard, its scatter plots in Figures 8 and  9 demonstrate that the resulting distributions $f_{\sigma^{2}}$ also have different behaviors between chaotic and nonchaotic cases. Their axis values indicate that they can be compared together in one plot. So we fit 333 The details of this data fitting is shown in the Appendix. their histogram to the normal distribution and find the average value $\langle\sigma^{2}\rangle$ and the standard deviation, which can characterize the resulting distribution $f_{\sigma^{2}}$. Then the resulting distribution $f_{\sigma^{2}}$ can be shown in one plot where the point is this average value and the error bar is the standard deviation. In Figure 5(a), the resulting distributions $f_{\sigma^{2}}$ are shown for all different distributions (GOE, GUE, URE, UIM and UCP) of initial operators, as the system varies from nonchaotic $a=0$ to chaotic $a=1$. When the system is nonchaotic, the data points of all distributions overlap together within the range of error bars. As the system becomes chaotic, different distributions start to separate approximately at $a=0.1$ and the distance of separation is roughly fixed after $a=0.15$. Note that the axis is logarithmic, so the separation is very large. We see that they split into two groups well-separated: one group is of GOE, URE and UIM, the other group is GUE and UCP.
For the Stadium billiard, the resulting distributions $f_{\sigma^{2}}$ always split into two well-separated groups as shown in Figure 6(a), whether the system is chaotic or not. Note that this is not a logarithmic plot because the values are nearly constants. Interestingly, the two groups match those observed in the case of the Sinai billiard of Figure  5(a). Comparing the values in these two plots, we see that the average $\langle\sigma^{2}\rangle$ is nearly identical: in one group it remains approximately $0.00035$ while in the other it stays around $0.00025$. From these values of the resulting distributions $f_{\sigma^{2}}$, the Stadium billiard exhibits behavior similar to the chaotic Sinai billiard.
5 Conclusion
In this paper, we studied the statistics involved by Lanczos coefficients in the Krylov operator complexity and deeply exhibited the connection among RMT, Anderson localization and Krylov complexity. We focused on the average correlation matrix $\langle x_{i}x_{j}\rangle$ and the resulting distribution $f_{\sigma^{2}}$ of the variance. We found in a phenomenological way that they satisfy the Wishart distribution and the (rescaled) chi-square distribution respectively, independent of the distributions of initial operators. With the help of these two quantities, we explored the Sinai and the Stadium billiard as a numerical example. The statistical distributions are verified and further interesting behaviors are obtained as the system changes from nonchaotic to chaotic, which indicates a consistent connection among RMT, Anderson localization and Krylov complexity.
For the Sinai billiard, the nonchaotic case is clearly distinguished from the chaotic case. The nonchaotic case has more erratic behavior as captured by $\left\langle x_{i}x_{j}\right\rangle$ which is consistent with the analysis of Dymarsky:2019elm ; Rabinovici:2021qqt , and has heavy cross-like structure as capture by $\ln{|\psi_{2m}\psi_{2n}|}$ which locally looks like two-electron correlations of Anderson localization Weinmann1995h2eOF . The resulting distribution $f_{\sigma^{2}}$ from different ensembles are overlapped with each other. However, for the Stadium billiard, there is no distinction between the nonchaotic and the chaotic case. More interestingly, the whole behavior of the Stadium billiard is similar to the chaotic Sinai billiard. Intuitively we can see that they have similar behavior in the chaotic case. But why the nonchaotic Stadium billiard is also similar to the chaotic case? The consistency among the three quantities $\left\langle x_{i}x_{j}\right\rangle$, $\ln{|\psi_{2m}\psi_{2n}|}$ and $f_{\sigma^{2}}$ indicates this result is reliable. So there should be a reason for it. Currently this is an open question.
@@ 5 Conclusion
Another interesting behavior is the two groups of separation for the resulting distribution $f_{\sigma^{2}}$. The GOE, URE and UIM are in one group. We can view GOE, URE and UIM as essentially composed of only real numbers, because the UIM is just real numbers multiplied by the overall factor-the imaginary $i$. The other group is GUE and UCP, which are essentially composed of complex numbers with both real and imaginary parts. This separation seems to be connected with the dimension of random numbers, which might be related to the nature of the integrability-breaking-term or to the symmetry of the model haake1991quantum . This remains to be one of the open problems.
In addition, we emphasize the importance of the resulting distribution being the normal distribution. We repeat the computation for $\mathcal{N}_{max}=5$ in Appendix, where the resulting distribution $f_{\sigma^{2}}$ is the rescaled chi square and not normal. Then there is no separation of $f_{\sigma^{2}}$ in all cases. Note that this rescaled chi square behavior of $\mathcal{N}_{max}=5$ further proves the phenomenological analysis in Section 3. Because few energy levels are kept with $\mathcal{N}_{max}=5$ it is impossible to extract the statistics numerically in a bottom-up manner. However, once the statistics is given by phenomenological analysis, the numerical results of $\mathcal{N}_{max}=5$ can prove the statistics in a top-down manner.
Qualitatively the resulting distribution being normal means that we have enough randomness in the sample, so that the statistics is reliable and robust. Empirically, a choice of $\mathcal{N}_{max}=50$ is already adequate to get the normal distribution. It remains an open problem to quantitatively understand the connection between being normal distributions and the separation behavior of $f_{\sigma^{2}}$.
@@
Acknowledgements.
Appendix A Numerical details not relevant to the main physics
Here we only show the numerical details for the Sinai billiard to keep the conciseness of the paper. For the Stadium billiard, most results are parallel except that there is no overlapping behavior in the nonchaotic case of $\mathcal{N}_{max}=100$. So there is no overlapping behavior in its plots analogous to Part [II] below, which has already be quantized in Figure 6(a) in the main context.
[I]: Firstly an example of Lanczos coefficients $b_{n}$ is shown in Figure 7, for the case of GUE. There is a clear difference between the chaotic and the nonchaotic case, which is explored further by the average correlation matrix $\left\langle x_{i}x_{j}\right\rangle$. The set of coefficients $\left\{b_{n}\right\}$ between the two red lines, $5\mathcal{N}_{max}\leq n\leq 10\mathcal{N}_{max}$, is selected to compute the variance $\sigma^{2}$. Note that this selection of the set does not qualitatively affect the result of this paper. The effect of selecting the set of $\left\{b_{n}\right\}$ from different algorithm step sizes is discussed in the end of this Appendix.
[II]: Then the variances $\sigma^{2}$ are shown in Figure 8 for the samples of different distributions (GOE, GUE, URE, UIM and UCP) of initial operators, where $m$ labels the sample. Note that only the first 1000 out of the total 5000 samples are displayed for visual clarity, as the remaining samples exhibit similar behavior. For GOE and GUE, they overlap completely in the nonchaotic case in Figure 8(b), and separate from each other in the chaotic case in Figure 8(a). Similary for the uniform distributions, they mix together in the nonchaotic case in Figure 8(d), and split into two separated groups in the chaotic case in Figure 8(c), with URE and UIM being one group and UCP being the other.
The histogram of all the sampled $\sigma^{2}$ is shown in Figure 9 and the resulting probability distribution resembles the normal distribution. For the Gaussian ensembles, we can see clearly that they are overlapped in the nonchaotic case in Figure 9(b) and separated in the chaotic case in Figure 9(a). For the uniform distributions, they are overlapped in the nonchaotic case in Figure 9(d). In the chaotic case in Figure 9(c), URE and UIM are still together, but they are separated from UCP.
Now we show the data fitting of the histogram in Figure 9 to the normal distribution. For GOE and GUE, the fitted distribution is shown in Figure 10. For uniform distributions, the fitted distribution is shown in Figure 11. The average value and the standard deviation are listed in Table 1, where we compare the values obtained from the sampled data and from the fitted normal distribution.
Note that the corresponding results of the Stadium billiard are always separating whether the system is chaotic or not, that it they are similar to Figure 8(a) and Figure 9(a).
[III]: After that, let’s verify the chi-square distribution of $\sigma^{2}$ for the case $\mathcal{N}_{max}=5$. The sets of coefficients $b_{n}$ are chosen from step-sizes $151<n<$ 15. We repeat the computation and show that there is neither normal distributions nor separating behaviors.
The histogram of all $5000$ samples are shown in Figure 12. Obviously they are not normal distributions and are not separated in the chaotic case. So the normal distribution is necessary for the separating behavior.
We fit the histogram to the rescaled chi-square distribution, the results are shown in Figure 13 for the Gaussian ensembles and in Figure 14 for the uniform distributions. We see that the rescaled chi-square distribution indeed captures the resulting distribution $f_{\sigma^{2}}$, so it verifies our analysis in Section 3.3.
Note that the corresponding results of the Stadium billiard are the same here, that it they are overlapping and satisfy the rescaled chi square distribution.
[IV]: Finally we show the independence on the sets of coefficients used in computing $\sigma^{2}$. Let’s choose the Lanczos coefficients $\left\{b_{n}\right\}$ from the new step-sizes $10\mathcal{N}_{max}\leq n\leq 15\mathcal{N}_{max}$ of Figure 7. Firstly, we compute the variance $\sigma^{2}$ (3.3), where the first $1000$ samples are shown in Figure 15. We see that they mix together in the nonchaotic case and separate from each other in the chaotic case. Then, the histogram of all $5000$ samples are shown in Figure 16. Obviously they still resemble normal distributions and are overlapped in the nonchaotic case and well-separated in the chaotic case. Finally, the fitted normal distributions are shown in Figure 17 and 18. So we still get the resulting normal distributions and the overlapping-splitting behaviors, although the detailed values of the average and the standard deviation change.
@@ Appendix A Numerical details not relevant to the main physics
(a) Histogram of $\sigma^{2}$ of GOE and GUE in the chaotic case $a=1$.
(b) Histogram of $\sigma^{2}$ of GOE and GUE in the nonchaotic case $a=0$.
(a) $N=5$ chaos, GUE. Skewness=1.19843, Kurtosis=5.18562.
(b) $N=5$ nonchaos, GUE. Skewness=4.30801, Kurtosis=50.7359.
(c) $N=5$ chaos, GOE. Skewness=1.6231, Kurtosis=7.05429.
(d) $N=5$ nonchaos, GOE. Skewness=5.83862, Kurtosis=73.6201.
(a) Chaotic case $a=1$ of UIM. Skewness=0.306957, Kurtosis=3.26486.
(b) Chaotic case $a=1$ of URE. Skewness=0.234259, Kurtosis=2.88038.
(c) Chaotic case $a=1$ of UCP. Skewness=0.207669, Kurtosis=2.97609.
x